The probability that A hits a target is 1/4, and the probability that B hits the target is 2/5. Both shoot at the target. What is the probability that at least one of them hits the target, i.e., that A or B (or both) hit the target?
Pick an option to check yourself, or open the answer below.
Show answer & explanation
Correct answer: D. 11/20
Explanation
P(at least one) = 1 − P(both miss) = 1 − (3/4)(3/5) = 1 − 9/20 = 11/20. (Equivalently 1/4 + 2/5 − 1/10 = 11/20.)More Discrete Structures and Optimisation questions
- If x and y are elements in a group G and if x⁵ = y³ = e, where e is the identity of G, then the inverse of x²yx⁴y² must be
- Arrange the following graphs on the basis of number of edges in increasing order (for large n): A. Kₙ (Complete Graph) B. Cₙ (Cycle graph)…
- A linear programming problem (LPP) is as follows: Min z = 30x − 18y, subject to the constraints 3x + 4y ≤ 60, 5x − 3y ≥ 20 and x, y ≥ 0. In…
- Which of the following graphs is/are planar? Graph A (vertices a–f): edges ab, ac, bd, cf, de, ef, bf, cd. Graph B (vertices a–f): edges af…
- Which of the following statements are true about sets? A. 0 ∈ ∅ B. ∅ ∈ {0} C. ∅ ∈ {∅} D. {∅} ∈ {∅} E. {∅} ⊂ {∅, {∅}} Choose the correct ans…
- Consider the following relations on {1, 2, 3, 4}. Which of the following relations are reflexive? (A) R₁ = {(1,1),(1,2),(2,1),(2,2),(3,4),(…